# Standardized Constraints on the Shadow Radius and the Instability of Scalar, Electromagnetic, $p$-Form, and Gravitational Perturbations of High-Dimensional Spherically Symmetric Black Holes in Einstein-power-Yang-Mills-Gauss-Bonnet Gravity

### v4

@@ V.3 Quasinormal frequencies and time-domain profiles
Since larger values of the Gauss-Bonnet coupling constant ${\alpha}_{2}$ can lead to dynamical instability, we aim to identify an interval for ${\alpha}_{2}$ that ensures dynamical stability. Since $\mathcal{Q}=0$ corresponds to the maximum domain of ${\alpha}_{2}$ values, we plotted the time-domain profiles of perturbations for the EGB black hole (19) with larger values of the parameter ${\alpha}_{2}$, as shown in the left panels of Figures 21 and 22. We further characterize the dynamical evolution through time-domain profiles of perturbed EYM black holes (20), as demonstrated in the right panels of Figures 21 and 22. We also presented the time-domain profiles corresponding to large $\ell$ values. Since the grid width in Eq. (206) is set to $\Delta=0.1$, in order to ensure the effectiveness of accuracy, we select $\ell=9$ as the display. Naturally, for larger values of $\mathcal{Q}$, perturbations also exhibit dynamical instability. However, since $\mathcal{Q}$ has only a minimal impact on the perturbations, we do not focus on identifying the stable region of the parameter $\mathcal{Q}$. Therefore, we concentrate exclusively on the allowable range of the Gauss-Bonnet coupling constant ${\alpha}_{2}$ in the context of high-dimensional EGB black hole solutions. By combining constraints from the high-dimensional shadow radius with the time-domain stability analysis of perturbations, we establish rigorously constrained parameter bounds, as comprehensively detailed in Table 4. Notably, the investigation into dynamical stability specifically targets gravitational perturbations (tensor modes and Regge-Wheeler modes) due to their greater physical relevance. Our findings reveal perfect concordance between the constraint intervals of the tensor and Regge-Wheeler modes, with the derived parameter bounds showing complete insensitivity to variations in the parameters $q$ and $\ell$. In Table 4, the parameters are set to $\mathsf{M}=2\pi\mathscr{R}=1$, G=1{G}=1 and $\mathscr{G}=1$.
@@ VI Conclusions
    Through the application of the high-dimensional shadow constraint formula (95) to EGBPYM black holes, we rigorously established that even in high-dimensional space-time frameworks, a physically viable domain of characteristic parameters remains definable. This result directly addresses and rectifies inconsistencies found in previous analyses of parameter bounds within this theoretical context. We observe that the value of the shadow radius $R_{s}$ will decrease as the coupling constant ${\alpha}_{2}$ increases, while the valid interval of the coupling constant ${\alpha}_{2}$ contracts monotonically with increasing magnetic charge $\mathcal{Q}$. Crucially, this valid interval will progressively shrink to fit within the admissible region $r_{\mathrm{sh}}$ defined by the high-dimensional shadow radius. In other words, for each space-time dimension, the valid range of the parameter ${\alpha}_{2}$ in EGB black holes ($\mathcal{Q}=0$) precisely matches the maximum permissible range of ${\alpha}_{2}$ in EGBPYM black holes. All in all, these results clearly indicate that under the constraints of the high-dimensional shadow constraint formula (95), the characteristic parameters ${\alpha}_{2}$ and $\mathcal{Q}$ exhibit well-defined valid intervals limited by observational data, demonstrating the dimensional dependence of parameter constraints in extended space-time geometries.
@@
    We provide a comprehensive display of the time-domain profiles for spin-0, spin-1, $p$-form, and spin-2 perturbations in various dimensions, from which we can intuitively obtain the following two significant features: First, the perturbations exhibit distinct parity behavior between odd and even dimensions, with the ring-down stage of the even-dimensional perturbations persisting longer than that of the odd-dimensional perturbations. Second, the influence of the power $q$ on the perturbations is minimal and can be regarded as negligible. It is only in the regime of large multipole numbers $\ell$ that the perturbation duration (ring-down phase) for larger values of $q$ exceeds that for smaller $q$. Furthermore, our analysis demonstrates that within EGBPYM space-time, spin-0, spin-1, $p$-form, and spin-2 perturbations exhibit dynamical stability in certain specific parameter regions across various spatiotemporal dimensions.

### v5

@@ V.3 Quasinormal frequencies and time-domain profiles
Since larger values of the Gauss-Bonnet coupling constant ${\alpha}_{2}$ can lead to dynamical instability, we aim to identify an interval for ${\alpha}_{2}$ that ensures dynamical stability. Since $\mathcal{Q}=0$ corresponds to the maximum domain of ${\alpha}_{2}$ values, we plotted the time-domain profiles of perturbations for the EGB black hole (20) with larger values of the parameter ${\alpha}_{2}$, as shown in the left panels of Figures 21 and 22. We further characterize the dynamical evolution through time-domain profiles of perturbed EYM black holes (21), as demonstrated in the right panels of Figures 21 and 22. We also presented the time-domain profiles corresponding to large $\ell$ values. Since the grid width in Eq. (204) is set to $\Delta=0.1$, in order to ensure the effectiveness of accuracy, we select $\ell=9$ as the display. Naturally, for larger values of $\mathcal{Q}$, perturbations also exhibit dynamical instability. However, since $\mathcal{Q}$ has only a minimal impact on the perturbations, we do not focus on identifying the stable region of the parameter $\mathcal{Q}$. Therefore, we concentrate exclusively on the allowable range of the Gauss-Bonnet coupling constant ${\alpha}_{2}$ in the context of high-dimensional EGB black hole solutions. By combining constraints from the high-dimensional shadow radius with the time-domain stability analysis of perturbations, we establish rigorously constrained parameter bounds, as comprehensively detailed in Table 4. Notably, the investigation into dynamical stability specifically targets gravitational perturbations (tensor modes and Regge-Wheeler modes) due to their greater physical relevance. Our findings reveal perfect concordance between the constraint intervals of the tensor and Regge-Wheeler modes, with the derived parameter bounds showing complete insensitivity to variations in the parameters $q$ and $\ell$. In Table 4, the parameters are set to $\mathsf{M}=M=1$ and $\mathscr{G}=1$.
@@ VI Conclusions
    Through the application of the high-dimensional shadow constraint formula (93) to EPYMGB black holes, we rigorously established that even in high-dimensional space-time frameworks, a physically viable domain of characteristic parameters remains definable. This result directly addresses and rectifies inconsistencies found in previous analyses of parameter bounds within this theoretical context. We observe that the value of the shadow radius $R_{s}$ will decrease as the coupling constant ${\alpha}_{2}$ increases, while the valid interval of the coupling constant ${\alpha}_{2}$ contracts monotonically with increasing magnetic charge $\mathcal{Q}$. Crucially, this valid interval will progressively shrink to fit within the admissible region $r_{\mathrm{sh}}$ defined by the high-dimensional shadow radius. In other words, for each space-time dimension, the valid range of the parameter ${\alpha}_{2}$ in EGB black holes ($\mathcal{Q}=0$) precisely matches the maximum permissible range of ${\alpha}_{2}$ in EPYMGB black holes. All in all, these results clearly indicate that under the constraints of the high-dimensional shadow constraint formula (93), the characteristic parameters ${\alpha}_{2}$ and $\mathcal{Q}$ exhibit well-defined valid intervals limited by observational data, demonstrating the dimensional dependence of parameter constraints in extended space-time geometries.
@@
    We provide a comprehensive display of the time-domain profiles for spin-0, spin-1, $p$-form, and spin-2 perturbations in various dimensions, from which we can intuitively obtain the following two significant features: First, the perturbations exhibit distinct parity behavior between odd and even dimensions, with the ring-down stage of the even-dimensional perturbations persisting longer than that of the odd-dimensional perturbations. Second, the influence of the power $q$ on the perturbations is minimal and can be regarded as negligible. It is only in the regime of large multipole numbers $\ell$ that the perturbation duration (ring-down phase) for larger values of $q$ exceeds that for smaller $q$. Furthermore, our analysis demonstrates that within EPYMGB space-time, spin-0, spin-1, $p$-form, and spin-2 perturbations exhibit dynamical stability in certain specific parameter regions across various spatiotemporal dimensions.